# ReCon: Enhancing True Correspondence Discrimination through Relation Consistency for Robust Noisy Correspondence Learning

### v1

@@ 1 Introduction
Figure 1: Illustration of our motivation. The relation-aware alignment correctly identifies the mismatched pair as negatives, while the relation-agnostic alignment fails to detect such inconsistency.
@@
As shown in Fig. 1(a), whether through relation-agnostic or relation-aware alignment, the true correspondence is expected to consistently assigned a high similarity score due to its perfect matching of both objects across modalities and relations within modalities. However, such matching is irreversibly compromised by the presence of untouchable noisy correspondence, thus narrows the distance between mismatches. Specifically, the unwanted misalignment erroneously reduces the distance of unassociated objects, inadvertently confusing retrieval models and thus undermining their discriminability for true correspondences. Besides, such misalignment also impairs the relations between objects within modalities, which significantly disrupts the contextual semantic consistency that is essential for true correspondences despite the nuance from objects. As shown in Fig. 1(b), the noisy pair with similar objects cannot be correctly identified by the relation-agnostic alignment due to its inability to effectively recognize the discrepancies of relations within modalities. Such misidentification inevitably introduces false supervisions, which misleads the model towards further wrong optimization direction. In contrast, the relation-aware alignment accurately identifies it as a negative pair, benefiting from its dual consideration of both cross-modal and intra-modal relations.
@@
Motivated by the above observations, we propose a general Relation Consistency learning framework, namely ReCon, to effectively mitigate the adverse impact caused by NCs, as shown in Fig. 2. The main motivation of ReCon is to enhance the discriminability for true correspondences. Specifically, a novel relation consistency learning is introduced to enable alignment not only between objects across modalities but relations within modalities. In details, the cross-modal relation consistency is presented to maximize the similarity scores of positive pairs while minimizing the negatives, ensuring that aligned objects have similar semantic representations. Meanwhile, the intra-modal relation consistency is designed to minimize the distance of relation matrices that describe the contextualized semantics of objects within modalities, which further enlarges the distinguish between positives and negatives. In practice, due to the lack of explicit annotations of objects, we propose to align the relation matrix extracted from one selected anchor modality with the proxy relation matrix extracted from another modality. Subsequently, such dual constraints of relations are employed to divide the noisy training data, wherein the divided partitions will be trained with corresponding strategies to achieve robust cross-modal retrieval, which significantly enhances the discriminability for true correspondences and effectively avoids the wrong supervisions of misidentified negatives.
@@
In summary, the main contributions of our work are as follows: (1) A general Relation Consistency learning framework, namely ReCon, is robustly proposed to identify the true correspondences and therefore mitigate the adverse impact caused by NCs within multimodal dataset. (2) A novel relation consistency learning approach is explicitly designed to jointly enforce the alignment of the cross-modal relation consistency and the intra-modal relation consistency. (3) A reliable true correspondence discrimination strategy is effectively presented to accurately partition the noisy data pairs, which therefore, seamlessly minimizes the risk caused by wrong supervisions and mitigates the misidentification of mismatches. (4) Extensive experiments highlight the advantages of the proposed ReCon in comparison with other SOTA methods and demonstrate its outstanding performances in challenging NCs scenario.
@@ 3.1 Preliminaries
Relation Consistency Learning
Given a sequence $\mathbf{O}=[o_{1},\cdots,o_{N_{o}}]$ containing $N_{o}$ objects appeared in a visual-text pair, the sequences of visual and linguistic can be denoted as $\overline{\mathbf{V}}=[\overline{v}_{1},\cdots,\overline{v}_{N_{o}}]$ and $\overline{\mathbf{L}}=[\overline{l}_{1},\cdots,\overline{l}_{N_{o}}]$, respectively. Here, each item with same index corresponds to a same object. Note that an object may be described by one or more words in the sentence and one or more regions in the image, such that the linguistic item and visual item may represent a collocation of words and regions, respectively. Moreover, the relation $\mathbf{C}_{o_{i}}=[c_{o_{i}\rightarrow o_{1}},\cdots,c_{o_{i}\rightarrow o_{N_{o}}}]$ of one object to others can also be depicted in both visual and linguistic modalities, which can be expressed as $\mathbf{C}_{\overline{v}_{i}}=[c_{\overline{v}_{i}\rightarrow\overline{v}_{1}},\cdots,c_{\overline{v}_{i}\rightarrow\overline{v}_{N_{o}}}]$ and $\mathbf{C}_{\overline{l}_{i}}=[c_{\overline{l}_{i}\rightarrow\overline{l}_{1}},\cdots,c_{\overline{l}_{i}\rightarrow\overline{l}_{N_{o}}}]$, respectively. Consequently, the relation consistency can be preserved by minimizing the expected risk for the semantic distance objective, as expressed in the following equation:
@@
where $\mathcal{L}_{SD}$ is the loss function that utilized for narrowing semantic distance, e.g., symmetric matrix-based Kullback-Leibler Divergence (m-LK) [29].
@@ 3.2 Relation Consistency Learning
Intra-modal relation consistency refers to the matching of semantics between visual contexts among regions and linguistic contexts among words. Unfortunately, the absence of explicit object annotations presents a particularly intricate and demanding challenge, which is quite common in real-world scenarios. Consequently, we cannot access to the sequences containing objects, which means that each visual/linguistic item now corresponds to only one region/word. Undoubtedly, Eq. (1) cannot be directly applied to such input sequences due to the lack of one-to-one correspondence properties found in object sequences. To address this problem, we first select an anchor modality, e.g., visual modality, containing regions sequence, and then construct a proxy sequence containing the most corresponding words sequence from the opposite modality, such that the relations of distinct modalities can be comparable. More specific, given a sequence $V$ with visual modality as anchor, the relations to sequence $L$ from the opposite modality can be obtained by $\mathbf{C}_{\mathcal{VL}}=g(V,L)$, wherein the relations between every visual item $v_{i}$ to all the linguistic items are depicted in $\mathbf{C}_{\mathcal{VL}}[i,:]$, i.e., the $i$-th row of this relation matrix. Subsequently, we can obtain the most relevant item $l_{i^{*}}$ for $v_{i}$, wherein the index can be calculated as $i^{*}=\arg\max\mathbf{C}_{\mathcal{VL}}[i,:]$. Likewise, we can obtain the most relevant linguistic item $l_{j^{*}}$ for the visual item $v_{j}$. Therefore, the intra-modal relations $c_{v_{i}\rightarrow v_{j}}$ within visual modality can be depicted by the intra-modal relations $c_{l_{i^{*}}\rightarrow l_{j^{*}}}$ within linguistic modality, which can be formulated in the following equation:
@@ 3.3 True Correspondence Discrimination
Inspired by the previous success [16, 31], we also leverage the small-loss criterion to achieve a rough division for the corrupted training data. Specifically, we first compute the per-sample cross-modal relation loss by $\mathcal{L}_{CM}$, denoted as $\{l_{i}^{CM}\}_{i=1}^{N}=\{\mathcal{L}_{CM}(V_{i},L_{i})\}_{i=1}^{N}$. Next, a two-component Gaussian Mixture Model (GMM) [25] would be employed to fit the per-sample loss distribution of all training pairs, which can be expressed as $p(l_{i}^{CM})=\sum_{k=1}^{K}\lambda_{k}\phi(l_{i}^{CM}|k)$. Here $K=2$, $\lambda_{k}$ is the corresponding mixture coefficient, and $\phi(l_{i}^{CM}|k)$ indicates the probability density function of the $k$-th component. Besides, the Expectation-Maximization [5] algorithm is employed to optimize the GMM. Finally, we use the component with smaller mean to obtain the estimated probability:
@@ 4.1 Datasets and Protocols
Recall at K (R@K) is a widely-used metric to measure the retrieval performance, defined as the percentage of matched items successfully retrieved from the top K candidates. In our experiments, the R@1, R@5, R@10, and the sum of three recalls for image-to-text and text-to-image retrieval are all reported to provide a comprehensive performance evaluation for our method.
@@ 4.2 Implementation Details
For fair comparisons, all experiments are conducted using the same backbone SGRAF [6] and all experimental settings are consistent with NCR [16], except for the specific parameters of ReCon. Specifically, the batch size $N_{b}$ is set to 128 and the temperature coefficients $\tau$ is 0.1. The division thresholds $\omega_{1}$ and $\omega_{2}$ are both set to 0.5, the scale parameter $\alpha$ for penalization factor is set to 0.1, and the momentum coefficient $\beta$ is 0.6. Moreover, the fixed margin $\gamma$ is set to 0.2 and the balance factor $\xi$ is 5. Before training models, we conduct a $\eta=5$ epochs warmup process for initial convergence. Besides, all experiments are conducted without any additional preprocessing or the use of external data sources. Due to the space limitation, more analysis and implementation details are given in our supplementary material.
@@ 5 Conclusion
This paper introduces a novel Relation Consistency learning framework, namely ReCon, to effectively mitigate the adverse impact caused by NCs. The main motivation of our ReCon is to enhance the discriminability of models for true correspondences in noisy multimodal dataset and thus effectively avoids the wrong supervisions of false correspondences, especially in the presence of hard NCs. Specifically, we leverage the dual constrains, which simultaneously consider the cross- and intra-modal relations, to jointly divide the corrupted training data into different partitions. Extensive experiments conducted on three widely-used cross-modal benchmarks validate the effectiveness and robustness of ReCon in handling both simulated and real-world NCs.

### v2

@@ 1 Introduction
Figure 1: Illustration of relation discrepancy. The relation-aware alignment correctly identifies mismatched pair as negatives, while relation-agnostic alignment fails to detect such inconsistency.
@@
As mentioned in IAIS work [29], the relation consistency often enhances the contextualized representation of image-text pairs. Inspired by this finding, we consider the relation discrepancy to mitigate the adverse impacts caused by mismatches among dataset. As shown in Fig. 1(a), whether through relation-agnostic or relation-aware alignment, the true correspondence is expected to consistently assigned a high similarity score due to its perfect matching of both objects across modalities and relations within modalities. However, such matching is irreversibly compromised by the presence of untouchable noisy correspondence, thus narrows the distance between mismatches. Specifically, the unwanted misalignment erroneously reduces the distance of unassociated objects, inadvertently confusing retrieval models and thus undermining their discriminability for true correspondences. Besides, such misalignment also impairs the relations between objects within modalities, which significantly disrupts the contextual semantic consistency that is essential for true correspondences despite the nuance from objects. As shown in Fig. 1(b), the noisy pair with similar objects cannot be correctly identified by the relation-agnostic alignment due to its inability to recognize the discrepancies of relations within modalities. Such misidentification inevitably introduces false supervisions, which misleads the model towards further wrong optimization direction. In contrast, the relation-aware alignment accurately identifies it as a negative pair, benefiting from its dual consideration of both cross-modal and intra-modal relations.
@@
Motivated by the above observations, we propose a general Relation Consistency learning framework, namely ReCon, to effectively mitigate the adverse impact caused by NCs, as shown in Fig. 2. The main motivation of ReCon is to enhance the discriminability for true correspondences. Specifically, an effective relation consistency alignment strategy is introduced to enable alignment not only between objects across modalities but relations within modalities. In details, the cross-modal relation consistency is presented to maximize the similarity scores of positive pairs while minimizing the negatives, ensuring that aligned objects have similar semantic representations. Meanwhile, the intra-modal relation consistency is employed to minimize the distance of relation matrices that describe the contextualized semantics of objects within modalities, which further enlarges the distinguish between positives and negatives. In practice, due to the lack of explicit annotations of objects, we propose to align the relation matrix extracted from one selected anchor modality with the proxy relation matrix extracted from another modality. Subsequently, such dual constraints of relations are employed to divide the noisy training data, wherein the divided partitions will be trained with corresponding strategies to achieve robust cross-modal retrieval, which significantly enhances the discriminability for true correspondences and effectively avoids the wrong supervisions of misidentified negatives.
@@
In summary, the main contributions of our work are as follows: (1) A general Relation Consistency learning framework, namely ReCon, is robustly proposed to identify the true correspondences and therefore mitigate the adverse impact caused by NCs within multimodal dataset. (2) An effective relation consistency alignment strategy is explicitly employed to jointly enforce the alignment of the cross-modal relation consistency and the intra-modal relation consistency. (3) A reliable true correspondence discrimination strategy is effectively presented to accurately partition the noisy data pairs, which therefore, seamlessly minimizes the risk caused by wrong supervisions and mitigates the misidentification of mismatches. (4) Extensive experiments highlight the advantages of the proposed ReCon in comparison with other SOTA methods and demonstrate its outstanding performances in challenging NCs scenario.
@@ 3.1 Preliminaries
Intra-Modal Relation Alignment
Given a sequence $\mathbf{O}=[o_{1},\cdots,o_{N_{o}}]$ containing $N_{o}$ objects appeared in a visual-text pair, the sequences of visual and linguistic can be denoted as $\overline{\mathbf{V}}=[\overline{v}_{1},\cdots,\overline{v}_{N_{o}}]$ and $\overline{\mathbf{L}}=[\overline{l}_{1},\cdots,\overline{l}_{N_{o}}]$, respectively. Here, each item with same index corresponds to a same object. Note that an object may be described by one or more words in the sentence and one or more regions in the image, such that the linguistic item and visual item may represent a collocation of words and regions, respectively. The relation $\mathbf{C}_{o_{i}}=[c_{o_{i}\rightarrow o_{1}},\cdots,c_{o_{i}\rightarrow o_{N_{o}}}]$ of one object to others can also be depicted in both visual and linguistic modalities, i.e., $\mathbf{C}_{\overline{v}_{i}}=[c_{\overline{v}_{i}\rightarrow\overline{v}_{1}},\cdots,c_{\overline{v}_{i}\rightarrow\overline{v}_{N_{o}}}]$ and $\mathbf{C}_{\overline{l}_{i}}=[c_{\overline{l}_{i}\rightarrow\overline{l}_{1}},\cdots,c_{\overline{l}_{i}\rightarrow\overline{l}_{N_{o}}}]$, respectively. Consequently, the alignment of such relations can be preserved by minimizing the expected risk for the distance objective [29], as expressed in the following equation:
@@
where $\mathcal{L}_{SD}$ is the loss function that utilized for narrowing semantic distance, e.g., symmetric matrix-based Kullback-Leibler Divergence (m-LK). Note that, IAIS [29] represents such relations within modalities using cross-modal attention matrix. Differently, ReCon obtains these relations by computing the similarity between objects within modalities.
@@ 3.2 Relation Consistency Learning
Intra-modal relation consistency refers to the matching of semantics between visual contexts among regions and linguistic contexts among words. Unfortunately, the absence of explicit object annotations presents a particularly intricate and demanding challenge, which is quite common in real-world scenarios. Consequently, we cannot access to the sequences containing objects, which means that each visual/linguistic item now corresponds to only one region/word. Undoubtedly, Eq. (1) cannot be directly applied to such input sequences due to the lack of one-to-one correspondence properties found in object sequences. To address this problem, like [29], we first select an anchor modality, e.g., visual modality, containing regions sequence, and then construct a proxy sequence containing the most corresponding words sequence from the opposite modality, such that the relations of distinct modalities can be comparable. Gven a sequence $V$ with visual modality as anchor, the relations to sequence $L$ from the opposite modality can be obtained by $\mathbf{C}_{\mathcal{VL}}=g(V,L)$, wherein the relations between every visual item $v_{i}$ to all the linguistic items are depicted in $\mathbf{C}_{\mathcal{VL}}[i,:]$, i.e., the $i$-th row of this relation matrix. Subsequently, we can obtain the most relevant item $l_{i^{*}}$ for $v_{i}$, wherein the index can be calculated as $i^{*}=\arg\max\mathbf{C}_{\mathcal{VL}}[i,:]$. Likewise, we can obtain the most relevant linguistic item $l_{j^{*}}$ for the visual item $v_{j}$. Therefore, the intra-modal relations $c_{v_{i}\rightarrow v_{j}}$ within visual modality can be depicted by the intra-modal relations $c_{l_{i^{*}}\rightarrow l_{j^{*}}}$ within linguistic modality, which can be formulated in the following equation:
@@ 3.3 True Correspondence Discrimination
Inspired by the previous success [15, 31], we also leverage the small-loss criterion to achieve a rough division for the corrupted training data. Specifically, we first compute the per-sample cross-modal relation loss by $\mathcal{L}_{CM}$, denoted as $\{l_{i}^{CM}\}_{i=1}^{N}=\{\mathcal{L}_{CM}(V_{i},L_{i})\}_{i=1}^{N}$. Next, a two-component Gaussian Mixture Model (GMM) [25] would be employed to fit the per-sample loss distribution of all training pairs, which can be expressed as $p(l_{i}^{CM})=\sum_{k=1}^{K}\lambda_{k}\phi(l_{i}^{CM}|k)$. Here $K=2$, $\lambda_{k}$ is the corresponding mixture coefficient, and $\phi(l_{i}^{CM}|k)$ indicates the probability density function of the $k$-th component. Besides, the Expectation-Maximization algorithm is employed to optimize the GMM. Finally, we use the component with smaller mean to obtain the estimated probability:
@@ 4.1 Datasets and Protocols
Recall at K (R@K) is a widely-used metric to measure the retrieval performance, defined as the percentage of matched items successfully retrieved from the top K candidates [22]. In our experiments, the R@1, R@5, R@10, and the sum of three recalls for image-to-text and text-to-image retrieval are all reported to provide a comprehensive performance evaluation for our method.
@@ 4.2 Implementation Details
For fair comparisons, all experiments are conducted using the same backbone SGRAF [5] and all experimental settings are consistent with NCR [15], except for the specific parameters of ReCon. Specifically, the batch size $N_{b}$ is set to 128 and the temperature coefficients $\tau$ is 0.1. The division thresholds $\omega_{1}$ and $\omega_{2}$ are both set to 0.5, the scale parameter $\alpha$ for penalization factor is set to 0.1, and the momentum coefficient $\beta$ is 0.6. Moreover, the fixed margin $\gamma$ is set to 0.2 and the balance factor $\xi$ is 5. Before training models, we conduct a $\eta=5$ epochs warmup process for initial convergence. Besides, all experiments are conducted without any additional preprocessing or the use of external data sources.
@@ 5 Conclusion
This paper introduces a general Relation Consistency learning framework, namely ReCon, to effectively mitigate the adverse impact caused by NCs. The main motivation of our ReCon is to enhance the discriminability of models for true correspondences in noisy multimodal dataset and thus effectively avoids the wrong supervisions of false correspondences, especially in the presence of hard NCs. Specifically, we leverage the dual constrains, which simultaneously consider the cross- and intra-modal relations, to jointly divide the corrupted training data into different partitions. Extensive experiments conducted on three widely-used cross-modal benchmarks validate the effectiveness and robustness of ReCon in handling both simulated and real-world NCs.